# The Symmetry Taco: Equivalences between Gapped, Gapless, and Mixed-State SPTs

### v1

@@ V.5 Symmetry Fractionalization and the Tetraptych Correspondence
Having introduced the SymTaco and illustrating that it captures the correct symmetry properties via the simple example of $\mathbb{Z}_{2}$ SWSSB, we now turn to a systematic analysis of gapped $1+1$d mixed-state SPTs with strong symmetry $K\lhd G$, where $G$ is the total symmetry group obtained via the group extension Eq. (51). Our goal in this Section is to explicitly derive the data characterizing the symmetry fractionalization of mixed-state SPTs (including iASPTs). We will find that this data exactly corresponds to that characterizing the pure-state folded SPTs discussed in Section IV. This is consistent with our argument that the valid set of anyon condensates for our mixed-state SymTFT is in one-to-one correspondence with the set of valid condensates for the pure-state SymTFT for folded states, in the case where the symmetry group is Abelian. We will also comment on the non-Abelian case and, doubling back to Sec. IV.2, argue that this correspondence allows us to indirectly obtain the 2-cocycle characterizing folded SPTs in the general non-Abelian case (which we were unable to directly derive in Section IV.2). Taken together, the results of this subsection complete the correspondences of phases in the tetraptych of Fig. 1.
Let us begin by deriving the symmetry fractionalization data of $1+1$d mixed-state SPTs (see also Refs. [74, 101, 100]). Recall from Section V.2 that we say that a mixed state is short-range entangled (and hence, in $1+1$d, an SPT) if its Choi state is short-range entangled. For our purposes, the essential consequences of assuming that $\rho$ is short-range entangled are that the strong and weak symmetries, when mapped to operators in the doubled space, fractionalize on the short-range entangled pure state |$\ket{\rho}\rangle$ [115]. Explicitly, if $U_{k}$ is the unitary representation of a strong symmetry $k\in K$, then its restriction to a sufficiently large interval $I$ on the short-range entangled state $\rho$ fractionalizes as
@@
where $U_{k}^{L,R}$ are unitary operators with support at the left and right endpoints of $I$. Similarly, under the assumption that |$\ket{\rho}\rangle$ is SRE in the Choi space, the restriction of a weak symmetry $g\in G$ implemented by $V_{g}$ to the region $I$ fractionalizes as
@@
where in the second equality, we used the fact that the combination $U_{g}^{L}U_{g}^{R}$ still forms a linear representation. Alternatively, we again fractionalize $U_{k}$ on $\rho$ but instead write
@@
On the other hand, we can repeat the above derivation in a slightly modified setting, where we act with a strong symmetry $k\in K$ on an interval $I$, followed by the action of an average symmetry $g\in G$ on on the entire system. We then find that
@@
We see that the symmetry fractionalization data $(\epsilon,\eta)$ of (intrinsically) ASPTs exactly parallels that of igSPTs reviewed in Sec. III.2. In an igSPT, the gapped degrees of freedom, charged under $A$, can form an $A$-SPT, while in an iASPT, the degrees of freedom charged under the strong symmetry $K$ can form an SPT; this data is captured by the 2-cocycle $\eta$. In the igSPT, the degrees of freedom charged under the low-energy symmetry group do not fractionalize as they are gapless while, in an iASPT, the weak symmetry simply cannot fractionalize. Finally, in an igSPT (iASPT), the fractionalized high-energy (strong) symmetry operators can carry charge under the low-energy (weak) symmetry; this data is captured by the function $\epsilon$. Repeating the same analysis as in Ref. [36], one can check that the pair $(\eta,\epsilon)$ characterizing an iASPT satisfy the same consistency conditions as those for an igSPT (see Eq. (19).) This correspondence suggests that it may be possible to generate iASPTs from igSPTs–for the $\mathbb{Z}_{4}$ iASPT, this process was discussed in Ref. [70]. We will return to this relation in Section VI from a SymTaco perspective and argue that a general protocol exists for generating iASPTs from igSPTs in $1+1$d.
From the SymTaco, we observe that the Choi state of an iASPT should also correspond to a $G\times G$ (folded) SPT. The 2-cocycle which characterizes this SPT in the doubled space is fully determined by $\eta$ and $\epsilon$. Defining $\omega((kg,g),(k^{\prime}g^{\prime},g^{\prime}))$ through
$U^{L}_{k}U^{L}_{g}U^{L}_{k^{\prime}}U^{L}_{g^{\prime}}\rho(U^{L}_{g}U^{L}_{g^{\prime}})^{\dagger}=\omega((kg,g),(k^{\prime}g^{\prime},g^{\prime}))$ (78)
×$\times U^{L}_{kgk^{\prime}g^{-1}}U^{L}_{gg^{\prime}}\rho U^{L\dagger}_{gg^{\prime}}\,,$ (79)
@@ V.6 Example: $\mathbb{Z}_{4}$ Intrinsically Average SPT
We can identify the $A$ and $B$ $\mathbb{Z}_{4}$ symmetries with the ket (+) and bra (-) $\mathbb{Z}_{4}$ symmetries, respectively. Note that, in the first line of Eq. (81), we are labeling the anyons as elements of $\mathcal{D}(G\times G)$, while the anyons in the second line are labeled as elements of $\mathcal{D}(G_{+})\boxtimes\overline{\mathcal{D}(G_{-})}$ where the -- anyons have spin opposite to the $+$ anyons. For instance, note that $e^{3}_{B}m_{B}$ has opposite braiding compared to $e_{A}m_{A}$. In the above set of anyons, we thus identify $e_{A}m_{A}$ with $e_{+}m_{+}$ and $e^{3}_{B}m_{B}$ with $e_{-}m_{-}$, so that $e_{A}m_{A}e_{B}m^{3}_{B}$ is identified with $e_{+}m_{+}\overline{e}_{-}\overline{m}_{-}$ in the second line.
@@ VI Applications of the Mixed-State SymTaco
In the preceding Section, we introduced the Choi state of the quantum double $\mathcal{D}(G)$–the SymTaco–as a mixed-state SymTFT for $1+1$d $G$-symmetric systems, where $G$ contains both strong and weak symmetries. Using the SymTaco, we established one of the main results of the present work: a one-to-one correspondence between intrinsically average SPTs, intrinsically gapless SPTs, and “folded” gapped SPTs. Now, beyond providing a formal tool for classifying states, the SymTFT formalism in the pure-state setting also sheds light on other kinematic properties and manipulations of $G$-symmetric systems. As discussed in Section II, this includes the anomalies and gauging of a symmetry. In this Section, as a step towards understanding the full implications of the SymTaco  on the properties of mixed states, we study the process of weak or classical gauging and, in doing so, uncover a novel mixed-state anomaly in strongly and/or weakly $\mathbb{Z}_{2}$ symmetric systems. As another application of the SymTaco, we also demonstrate how our correspondence between igSPTs and iASPTs can be used to motivate explicit decoherence channels to generate the latter from the former. Finally, for Abelian symmetries $G$, we also develop a description of mixed-state SPTs in terms of Pauli subsystem codes, which may be of independent interest.
@@ VI.1 Gauging Weak Symmetries and a New Mixed-State “Anomaly”
As first observed in Ref. [72], the refined symmetry structure of mixed states opens the possibility for classically or weakly gauging a symmetry, i.e., gauging a weak symmetry. In brief, when gauging a symmetry in a pure state |$\ket{\psi}$, one projects the state to the subspace satisfying a certain Gauss’ law (see Refs [117, 118] for a discussion). For instance, for a spin-$1/2$ chain with spin-flip symmetry $X\equiv\prod_{i}X_{i}$, one adds ancilla degrees of freedom (the gauge fields) and enforces the Gauss’ law $G_{i}=Z_{i-1/2}X_{i}Z_{i+1/2}\equiv 1$, where the original qubits live on integer valued sites and the gauge fields live on the half-integer valued sites. The gauged state, |$\ket{\tilde{\psi}}$ satisfies $G_{i}\ket{\tilde{\psi}}=\ket{\tilde{\psi}}$ for all $i$. Gauging promotes the symmetry generated by $X$ to a local symmetry, in the sense that we can apply spin-flips to a finite region, at the expense of adding the gauge degrees of freedom. In contrast, when we classically gauge a symmetry of a mixed state $\rho$, we only demand that the gauged state, $\tilde{\rho}$, is invariant under conjugation by the Gauss’ law: $G_{i}\tilde{\rho}G_{i}^{\dagger}=\tilde{\rho}$. Physically, classical gauging may be understood as incoherently proliferating symmetry defects of the symmetry being weakly gauged.
At the level of states, this process may be understood in the doubled Hilbert space via the Choi isomorphism. Suppose $\rho$ is strongly symmetric under $K$, such that $U_{g}\rho=e^{i\theta_{g}}\rho$ for $g\in K$. Then, the vectorized density matrix |$\ket{\rho}\rangle$ is symmetric under $K\times K$. Classical or weak gauging then corresponds to applying the usual gauging procedure to the diagonal subgroup $K^{\mathrm{diag}}\subset K\times K$, rather than the full symmetry group $K\times K$ (which would describe the familiar “strong” gauging). This makes clear how we should implement classical gauging at the level of the SymTaco–one simply needs to exchange the electric reference boundary with that corresponding to spontaneous breaking down to $K^{\mathrm{diag}}$ which, in the original Hilbert space, describes strong-to-weak spontaneous symmetry breaking of the strong symmetry $K$. Note that, if the system only has a weak symmetry $G_{weak}$ to begin with, classically gauging it would correspond to exchanging the charge condensed boundary with that corresponding to weak-to-nothing breaking of the $G_{weak}$ symmetry, i.e., replacing the charge-condensed boundary with the original symmetry preserving boundary.
To illustrate this procedure, let us consider the example of $1+1$d states invariant under a strong $\mathbb{Z}_{2}$ symmetry, as discussed in Section V.4. Recall that the corresponding bulk TO in the Choi space is $\mathcal{D}(\mathbb{Z}_{2}^{+}\times\mathbb{Z}_{2}^{-})$, a bilayer Toric Code. The reference boundary of the SymTaco  is chosen to be the charge condensed boundary, with the corresponding Lagrangian subgroup given by $\mathcal{L}_{e}=\langle e_{+},e_{-}\rangle$. With this choice, the strong $\mathbb{Z}_{2}$ symmetry is generated by the magnetic anyons, $m_{+}$ and $m_{-}$. Now, to classically gauge the $\mathbb{Z}_{2}$ symmetry within the SymTaco, we exchange the charge condensed Lagrangian subgroup $\mathcal{L}_{e}$ on the reference boundary with the subgroup corresponding to SWSSB of the $\mathbb{Z}_{2}$ symmetry, namely to the Lagrangian subgroup, $\mathcal{L}_{1}=\langle e_{+}e_{-},m_{+}m_{-}\rangle$.
@@ VI.2 Fermionic SymTacos
Having built some intuition for bosonic systems with $\mathbb{Z}_{2}$ symmetry, let us briefly consider fermionic systems with only a weak $\mathbb{Z}_{2}^{F}$ fermion parity symmetry. In the pure-state setting, the SymTFT for fermionic systems has recently been developed in Refs. [44, 45, 43]. For a $1+1$d pure state with only $\mathbb{Z}_{2}^{F}$ fermion parity symmetry, the relevant SymTFT is the usual $2+1$d Toric Code with the reference boundary taken to be $\mathcal{L}_{f}=\langle 1,f\rangle$; here, the emergent fermion $f$ is understood to be condensed by adding an ancilliary fermionic chain (only along the reference boundary and not the dynamical boundary or the bulk) and pair condensing the $f$ anyons with the physical fermions. The ancilla fermions along the reference boundary encode the fact that the SymTFT describes a system composed of microscopic fermions. On the dynamical boundary however, we are still only permitted to condense bosonic anyons since, on physical grounds, only bosonic operators can condense. We can choose the fermion parity symmetry to be generated by the confined $m$ anyon. The two choices of gapped boundary for the dynamical boundary–the $e$ and $m$ condensates–then correspond to the same phases as the Kitaev chain and a trivially gapped chain, respectively. We can also gauge the fermion parity symmetry (which corresponds to bosonizing the system) by exchanging the $\mathcal{L}_{f}$ reference boundary with $\mathcal{L}_{e}$; this yields a bosonic system with $\mathbb{Z}_{2}$ symmetry generated by the $m$ line. In particular, the Kitaev chain and trivial phases, corresponding to the $e$ and $m$ condensed dynamical boundaries, map to the $\mathbb{Z}_{2}$-SSB and trivial symmetric phases, as expected under the familiar Jordan-Wigner transformation.
@@
Let us now consider a mixed state with only a weak $\mathbb{Z}_{2}^{F}$ fermion parity symmetry. We propose that the relevant SymTFT for such a system is provided by the $f$-decohered $2+1$d Toric Code discussed in Refs. [122, 71, 72]. In the Choi-state representation, this corresponds to starting with a bilayer Toric Code and condensing $f_{+}f_{-}$, such that the bulk anyons are generated by $\langle f_{+}\sim f_{-},e_{+}e_{-}\sim m_{+}m_{-}\rangle$. On the reference boundary, we then condense $\mathcal{L}_{f_{\pm}}=\langle f_{+}\sim f_{-}\rangle$ by introducing ancilla fermions. This leaves a single confined anyon on the boundary, $e_{+}e_{-}\sim m_{+}m_{-}$, which we take to be the generator of the weak $\mathbb{Z}_{2}^{F}$ fermion parity symmetry. This leaves only a single choice of Lagrangian subgroup for the dynamical boundary that respects positivity (recall that this must be bosonic): condensing $e_{+}e_{-}\sim m_{+}m_{-}$. This choice reflects the fact that, in the presence of a weak $\mathbb{Z}_{2}^{F}$ symmetry, there is only a single short-range-entangled phase given by the maximally mixed state. In particular, the distinction between the trivial phase and the Kitaev chain phase is removed when the strong $\mathbb{Z}_{2}^{F}$ is broken to the weak $\mathbb{Z}_{2}^{F}$ [123].
@@ VI.3 Pauli Subsystem Codes for Abelian Mixed-State SPTs
A natural setting in which nontrivial mixed states arise is in the context of quantum error-correcting codes, which are engineered to stabilize quantum information against environmental noise (up to certain thresholds). Indeed, there is a deep conceptual connection between mixed-state phases of matter and quantum error correction, since any resource-theoretic property of a specific mixed state is expected to persist throughout the entire mixed-state phase [111]. For topologically ordered phases in $2+1$d, the correspondence with quantum error correction is particularly well established. In the pure-state setting, for example, any $2+1$d Abelian anyon theory that admits a gapped boundary to vacuum can be fully described in terms of Pauli stabilizer codes [124]. More recently, it has been shown that any Abelian anyon theory (even non-modular theories) can be captured by topological subsystem codes [124] and that these are in correspondence with mixed-state topological orders [71, 72]. Motivated by these developments, we use the intuition provided by the SymTaco construction to develop a Pauli subsystem code description of mixed-state SPTs. Indeed, in the same way that the boundary anyon condensates in a pure-state SymTFT motivate explicit Hamiltonians, we find that the condensates in the SymTaco naturally correspond to subsystem codes. This yields an explicit form for the density matrices of $1+1$d mixed-state SPTs, from which their universal data can be extracted. Moreover, this approach naturally suggests a protocol for preparing $1+1$d mixed-state SPTs with Abelian symmetries, which we briefly discuss in Sec. VI.4.
@@
This channel may physically be understood as the process of measuring all the gauge operators and then forgetting the measurement outcomes. This results in a state that is maximally mixed within the gauge subsystem. Now, let |$\ket{\psi}$ be an arbitrary state in the code space. The TSSC code naturally describes the mixed state
$\rho_{\mathcal{G}}\equiv\mathcal{E}_{\mathcal{G}}[\ket{\psi}\bra{\psi}]\,.$ (90)
@@
We see that the stabilizer and gauge groups are intimately related to the strong and weak symmetries of a mixed state; precisely this structure was exploited in Refs. [71, 72] to show that $2+1$d TSSCs can be used to describe mixed-state topological orders. Note that if the logical space is empty, we can write
@@
We illustrate this subsystem code formalism for $1+1$d mixed-state phases discussed above using three examples of increasing complexity: (i) the SWSSB state, (ii) the decohered cluster state, and (iii) the $\mathbb{Z}_{4}$ intrinsically mixed SPT state.
@@
This is precisely the state studied in Ref. [69, 70]. Let us see how the symmetry leads to a symmetry-protected degeneracy when the system is placed on open boundary conditions. For a chain with $L$ sites labeled $10\ldots L-$ 1 with open boundary conditions, the symmetric stabilizer group is
@@
The mixed-state SPT, obtained via the decoherence procedure provided earlier in this Section, is the maximally mixed state within the subspace where all of the stabilizers are $+1$. This is given by the density matrix
@@
which is the projector onto the $+1$ eigenspace of all of the stabilizers. This density matrix was previously introduced in Sec. V.6, where we also saw how the universal data of symmetry fractionalization was realized on the density matrix $\rho_{\mathcal{G}}$. Here, we us explore a different feature of this intrinsically mixed SPT: its properties on a system with open boundary conditions.
Suppose the model is defined on a chain of length $L+1$, with sites ranging from 00 to $L$. The gauge operators are
@@
for $i$ ranging from 00 to $L-1$. We have only included gauge operators from Eq. (102) whose support is completely contained within 00 and $L$. Due to the removal of certain generators, the center becomes
@@
for $i$ ranging from 00 to $L-1$. Compared to the case of periodic boundary conditions, the center contains the Pauli operators $\mathcal{Z}_{0}^{2}\mathcal{X}_{0}^{2}$ and $\mathcal{Z}_{L}^{2}$, which are localized at the boundaries and are not symmetric under the total $\mathbb{Z}_{4}$ symmetry.
@@
corresponding to ±$\pm 1$ eigenvalues of the non-symmetric stabilizers $\mathcal{Z}_{0}^{2}\mathcal{X}_{0}^{2}$ and $\mathcal{Z}_{L}^{2}$.

### v2

@@ V.5 Symmetry Fractionalization and the Tetraptych Correspondence
Having introduced the SymTaco and illustrating that it captures the correct symmetry properties via the simple example of $\mathbb{Z}_{2}$ SWSSB, we now turn to a systematic analysis of gapped 1+11+1d mixed-state SPTs with strong symmetry $K\lhd G$, where $G$ is the total symmetry group obtained via the group extension Eq. (51). Our goal in this Section is to explicitly derive the data characterizing the symmetry fractionalization of mixed-state SPTs (including iASPTs). We will find that this data exactly corresponds to that characterizing the pure-state folded SPTs discussed in Section IV. This is consistent with our argument that the valid set of anyon condensates for our mixed-state SymTFT is in one-to-one correspondence with the set of valid condensates for the pure-state SymTFT for folded states, in the case where the symmetry group is Abelian. We will also comment on the non-Abelian case and, doubling back to Sec. IV.2, argue that this correspondence allows us to indirectly obtain the 2-cocycle characterizing folded SPTs in the general non-Abelian case (which we were unable to directly derive in Section IV.2). Taken together, the results of this subsection complete the correspondences of phases in the tetraptych of Fig. 1.
[…]
@@
[…]
@@
where in the second equality, we used the fact that the combination $U_{g}^{L}U_{g}^{R}$ still forms a linear representation. Alternatively, we again fractionalize $U_{k}$ on $\rho$ but instead write,
@@
On the other hand, we can repeat the above derivation in a slightly modified setting, where we act with a strong symmetry $k\in K$ on an interval II, followed by the action of an average symmetry $g\in G$ on on the entire system. We then find that,
@@
We see that the symmetry fractionalization data $(\epsilon,\eta)$ of (intrinsically) ASPTs exactly parallels that of igSPTs reviewed in Sec. III.2. In an igSPT, the gapped degrees of freedom, charged under $A$, can form an $A$-SPT, while in an iASPT, the degrees of freedom charged under the strong symmetry $K$ can form an SPT; this data is captured by the 2-cocycle $\eta$. In the igSPT, the degrees of freedom charged under the low-energy symmetry group do not fractionalize as they are gapless while, in an iASPT, the weak symmetry simply cannot fractionalize. Finally, in an igSPT (iASPT), the fractionalized high-energy (strong) symmetry operators can carry charge under the low-energy (weak) symmetry; this data is captured by the function $\epsilon$. Repeating the same analysis as in Ref. [37], one can check that the pair $(\eta,\epsilon)$ characterizing an iASPT satisfy the same consistency conditions as those for an igSPT (see Eq. (19).) This correspondence suggests that it may be possible to generate iASPTs from igSPTs–for the $\mathbb{Z}_{4}$ iASPT, this process was discussed in Ref. [73]. We will return to this relation in Section VI from a SymTaco perspective and argue that a general protocol exists for generating iASPTs from igSPTs in 1+11+1d.
From the SymTaco, we observe that the Choi state of an iASPT should also correspond to a $G\times G$ (folded) SPT. The 2-cocycle which characterizes this SPT in the doubled space is fully determined by $\eta$ and $\epsilon$. Defining $\omega((kg,g),(k^{\prime}g^{\prime},g^{\prime}))$ through,
[…]
@@ V.6 Example: $\mathbb{Z}_{4}$ Intrinsically Average SPT
We can identify the $A$ and $B$ $\mathbb{Z}_{4}$ symmetries with the ket (+) and bra (-) $\mathbb{Z}_{4}$ symmetries, respectively. Note that, in the first line of Eq. (80), we are labeling the anyons as elements of $\mathcal{D}(G\times G)$, while the anyons in the second line are labeled as elements of $\mathcal{D}(G_{+})\boxtimes\overline{\mathcal{D}(G_{-})}$ where the - anyons have spin opposite to the ++ anyons. For instance, note that $e^{3}_{B}m_{B}$ has opposite braiding compared to $e_{A}m_{A}$. In the above set of anyons, we thus identify $e_{A}m_{A}$ with $e_{+}m_{+}$ and $e^{3}_{B}m_{B}$ with $e_{-}m_{-}$, so that $e_{A}m_{A}e_{B}m^{3}_{B}$ is identified with $e_{+}m_{+}\overline{e}_{-}\overline{m}_{-}$ in the second line.
@@ VI Applications of the Mixed-State SymTaco
In the preceding Section, we introduced the Choi state of the quantum double $\mathcal{D}(G)$–the SymTaco–as a mixed-state SymTFT for 1+11+1d $G$-symmetric systems, where $G$ contains both strong and weak symmetries. Using the SymTaco, we established one of the main results of the present work: a one-to-one correspondence between intrinsically average SPTs, intrinsically gapless SPTs, and “folded” gapped SPTs. Now, beyond providing a formal tool for classifying states, the SymTFT formalism in the pure-state setting also sheds light on other kinematic properties and manipulations of $G$-symmetric systems. As discussed in Section II, this includes the anomalies and gauging of a symmetry. In this Section, as a step towards understanding the full implications of the SymTaco  on the properties of mixed states, we study the process of weak or classical gauging and, in doing so, uncover a novel mixed-state anomaly in strongly and/or weakly $\mathbb{Z}_{2}$ symmetric systems. As another application of the SymTaco, we also demonstrate how our correspondence between igSPTs and iASPTs can be used to motivate explicit decoherence channels to generate the latter from the former. Finally, for Abelian symmetries $G$, we also develop a description of mixed-state SPTs in terms of Pauli subsystem codes, which may be of independent interest.
@@ VI.1 Gauging Weak Symmetries and a New Mixed-State “Anomaly”
[…]
To illustrate this procedure, let us consider the example of 1+11+1d states invariant under a strong $\mathbb{Z}_{2}$ symmetry, as discussed in Section V.4. Recall that the corresponding bulk TO in the Choi space is $\mathcal{D}(\mathbb{Z}_{2}^{+}\times\mathbb{Z}_{2}^{-})$, a bilayer Toric Code. The reference boundary of the SymTaco  is chosen to be the charge condensed boundary, with the corresponding Lagrangian subgroup given by $\mathcal{L}_{e}=\langle e_{+},e_{-}\rangle$. With this choice, the strong $\mathbb{Z}_{2}$ symmetry is generated by the magnetic anyons, $m_{+}$ and $m_{-}$. Now, to classically gauge the $\mathbb{Z}_{2}$ symmetry within the SymTaco, we exchange the charge condensed Lagrangian subgroup $\mathcal{L}_{e}$ on the reference boundary with the subgroup corresponding to SWSSB of the $\mathbb{Z}_{2}$ symmetry, namely to the Lagrangian subgroup, $\mathcal{L}_{1}=\langle e_{+}e_{-},m_{+}m_{-}\rangle$.
@@ VI.2 Fermionic SymTacos
Having built some intuition for bosonic systems with $\mathbb{Z}_{2}$ symmetry, let us briefly consider fermionic systems with only a weak $\mathbb{Z}_{2}^{F}$ fermion parity symmetry. In the pure-state setting, the SymTFT for fermionic systems has recently been developed in Refs. [45, 46, 44]. For a 1+11+1d pure state with only $\mathbb{Z}_{2}^{F}$ fermion parity symmetry, the relevant SymTFT is the usual 2+12+1d Toric Code with the reference boundary taken to be $\mathcal{L}_{f}=\langle 1,f\rangle$; here, the emergent fermion $f$ is understood to be condensed by adding an ancilliary fermionic chain (only along the reference boundary and not the dynamical boundary or the bulk) and pair condensing the $f$ anyons with the physical fermions. The ancilla fermions along the reference boundary encode the fact that the SymTFT describes a system composed of microscopic fermions. On the dynamical boundary however, we are still only permitted to condense bosonic anyons since, on physical grounds, only bosonic operators can condense. We can choose the fermion parity symmetry to be generated by the confined $m$ anyon. The two choices of gapped boundary for the dynamical boundary–the $e$ and $m$ condensates–then correspond to the same phases as the Kitaev chain and a trivially gapped chain, respectively. We can also gauge the fermion parity symmetry (which corresponds to bosonizing the system) by exchanging the $\mathcal{L}_{f}$ reference boundary with $\mathcal{L}_{e}$; this yields a bosonic system with $\mathbb{Z}_{2}$ symmetry generated by the $m$ line. In particular, the Kitaev chain and trivial phases, corresponding to the $e$ and $m$ condensed dynamical boundaries, map to the $\mathbb{Z}_{2}$-SSB and trivial symmetric phases, as expected under the familiar Jordan-Wigner transformation.
@@
Let us now consider a mixed state with only a weak $\mathbb{Z}_{2}^{F}$ fermion parity symmetry. We propose that the relevant SymTFT for such a system is provided by the $f$-decohered 2+12+1d Toric Code discussed in Refs. [126, 74, 75]. In the Choi-state representation, this corresponds to starting with a bilayer Toric Code and condensing $f_{+}f_{-}$, such that the bulk anyons are generated by $\langle f_{+}\sim f_{-},e_{+}e_{-}\sim m_{+}m_{-}\rangle$. On the reference boundary, we then condense $\mathcal{L}_{f_{\pm}}=\langle f_{+}\sim f_{-}\rangle$ by introducing ancilla fermions. This leaves a single confined anyon on the boundary, $e_{+}e_{-}\sim m_{+}m_{-}$, which we take to be the generator of the weak $\mathbb{Z}_{2}^{F}$ fermion parity symmetry. This leaves only a single choice of Lagrangian subgroup for the dynamical boundary that respects positivity (recall that this must be bosonic): condensing $e_{+}e_{-}\sim m_{+}m_{-}$. This choice reflects the fact that, in the presence of a weak $\mathbb{Z}_{2}^{F}$ symmetry, there is only a single short-range-entangled phase given by the maximally mixed state. In particular, the distinction between the trivial phase and the Kitaev chain phase is removed when the strong $\mathbb{Z}_{2}^{F}$ is broken to the weak $\mathbb{Z}_{2}^{F}$ [127].
@@ VI.3 Pauli Subsystem Codes for Abelian Mixed-State SPTs
A natural setting in which nontrivial mixed states arise is in the context of quantum error-correcting codes, which are engineered to stabilize quantum information against environmental noise (up to certain thresholds). Indeed, there is a deep conceptual connection between mixed-state phases of matter and quantum error correction, since any resource-theoretic property of a specific mixed state is expected to persist throughout the entire mixed-state phase [115]. For topologically ordered phases in 2+12+1d, the correspondence with quantum error correction is particularly well established. In the pure-state setting, for example, any 2+12+1d Abelian anyon theory that admits a gapped boundary to vacuum can be fully described in terms of Pauli stabilizer codes [128]. More recently, it has been shown that any Abelian anyon theory (even non-modular theories) can be captured by topological subsystem codes [128] and that these are in correspondence with mixed-state topological orders [74, 75]. Motivated by these developments, we use the intuition provided by the SymTaco construction to develop a Pauli subsystem code description of mixed-state SPTs. Indeed, in the same way that the boundary anyon condensates in a pure-state SymTFT motivate explicit Hamiltonians, we find that the condensates in the SymTaco naturally correspond to subsystem codes. This yields an explicit form for the density matrices of 1+11+1d mixed-state SPTs, from which their universal data can be extracted. Moreover, this approach naturally suggests a protocol for preparing 1+11+1d mixed-state SPTs with Abelian symmetries, which we briefly discuss in Sec. VI.4.
@@
[…]
$|].\displaystyle\rho_{\mathcal{G}}\equiv\mathcal{E}_{\mathcal{G}}[\ket{\psi}\bra{\psi}$]\,. (89)
@@
We see that the stabilizer and gauge groups are intimately related to the strong and weak symmetries of a mixed state; precisely this structure was exploited in Refs. [74, 75] to show that 2+12+1d TSSCs can be used to describe mixed-state topological orders. Note that if the logical space is empty, we can write
@@
We illustrate this subsystem code formalism for 1+11+1d mixed-state phases discussed above using three examples of increasing complexity: (i) the SWSSB state, (ii) the decohered cluster state, and (iii) the $\mathbb{Z}_{4}$ intrinsically mixed SPT state.
@@
This is precisely the state studied in Ref. [72, 73]. Let us see how the symmetry leads to a symmetry-protected degeneracy when the system is placed on open boundary conditions. For a chain with $L$ sites labeled $0\ldots L-1$ with open boundary conditions, the symmetric stabilizer group is
@@
The mixed-state SPT, obtained via the decoherence procedure provided earlier in this Section, is the maximally mixed state within the subspace where all of the stabilizers are +1+1. This is given by the density matrix
@@
which is the projector onto the +1+1 eigenspace of all of the stabilizers. This density matrix was previously introduced in Sec. V.6, where we also saw how the universal data of symmetry fractionalization was realized on the density matrix $\rho_{\mathcal{G}}$. Here, we us explore a different feature of this intrinsically mixed SPT: its properties on a system with open boundary conditions.
Suppose the model is defined on a chain of length $L+1$, with sites ranging from 0 to $L$. The gauge operators are
@@
for $i$ ranging from 0 to L1L-1. We have only included gauge operators from Eq. (101) whose support is completely contained within 0 and $L$. Due to the removal of certain generators, the center becomes
@@
for $i$ ranging from 0 to L1L-1. Compared to the case of periodic boundary conditions, the center contains the Pauli operators $\mathcal{Z}_{0}^{2}\mathcal{X}_{0}^{2}$ and $\mathcal{Z}_{L}^{2}$, which are localized at the boundaries and are not symmetric under the total $\mathbb{Z}_{4}$ symmetry.
@@
corresponding to $\pm 1$ eigenvalues of the non-symmetric stabilizers $\mathcal{Z}_{0}^{2}\mathcal{X}_{0}^{2}$ and $\mathcal{Z}_{L}^{2}$.